# PUB: A Plasma-Propelled Ultra-Quiet Blimp with Two-DOF Vector Thrusting

### v1

@@ 2.1.1 Blimp Design
As shown in Figure 1(a), the main body of PUB is an ellipsoidal aluminum-film helium balloon of length $1.97\text{\,}\mathrm{m}$ and maximum diameter $51\text{\,}\mathrm{c}\mathrm{m}$. The volume is approximated by the ellipsoid formula
@@
The thruster layout has undergone several iterations. In the first-generation design (Figure 1(a)), six identical thrusters are rigidly attached at different positions, requiring simultaneous control of their on/off states and thrust magnitudes to realize six-DOF motion. This resulted in high system mass and control complexity. The second-generation design (Figure 1(b)) simplifies the layout to a single thruster mounted on a two-DOF platform directly beneath the envelope, enabling thrust vectoring in pitch and yaw and reducing weight while enhancing controllability.
@@ 2.1.2 Two-DOF Gimbal
As shown in Figure. 2, the gimbal uses two RC servos (each $9\text{\,}\mathrm{g}$; total gimbal mass $\approx$33 $3\text{\,}\mathrm{g}$; rated torque $1.6\text{\,}\mathrm{k}\mathrm{g}\,\mathrm{c}\mathrm{m}$) to vector the thruster. The upper stage rotates in the blimp’s oxzoxz symmetry plane to control pitch, with an angular range of $0^{\circ}$$180^{\circ}$. The lower stage is rigidly attached beneath the blimp and rotates about the $z$ axis to control yaw, also within $0^{\circ}$$180^{\circ}$.
@@ 2.2 Plasma Thruster
Figure 3 shows the thruster: a four-layer concentric ring frame with $9.44\text{\,}\mathrm{m}\mathrm{m}$ spacing between adjacent rings. The positive electrode is $0.1\text{\,}\mathrm{m}\mathrm{m}$ copper wire; the negative electrode is $40\text{\,}\mathrm{m}\mathrm{m}$-wide aluminum foil. The electrode separation is $25\text{\,}\mathrm{m}\mathrm{m}$.
@@ 3.1.2 Principle of Ionic-Wind Propulsion
The effect states that a properly shaped electrode pair immersed in a dielectric and biased at high voltage generates a net force (Figure 5).
@@
Based on the ion wind effect model from section 3.1.1, we can further simplify and form a thrust model for the ion propulsion device [12] [13]. Figure 6 illustrates the thrust direction.
@@ 3.2.1 Thrust Measurement
According to the plasma propulsion principles in section 3.1.2, the distance between the positive copper wire and the negative aluminum foil affects the thrust. To find the optimal spacing, we repeatedly changed the gap of the dual-ring thruster and measured the thrust. The experimental process was as follows: The thruster was inverted on a high-precision electronic scale ($0.01\text{\,}\mathrm{g}$ accuracy) with the negative electrode facing up, ensuring it was well-insulated from the scale. After taring the scale, the reading displayed the maximum thrust in grams when the power was turned on. The dual-ring thruster had a net weight of $16.00\text{\,}\mathrm{g}$. We conducted five parallel experiments, decreasing the electrode spacing each time. The thrust measurements are shown in Figure 7. We found that as the spacing decreased, the thrust gradually increased, reaching a maximum thrust-to-weight ratio of $0.7105\text{\,}\mathrm{N}\text{\,}{\mathrm{kg}}^{-1}$. However, when the gap was reduced to $2.5\text{\,}\mathrm{cm}$, a tip discharge phenomenon occurred, causing the negative aluminum foil to be punctured and burned. This demonstrated that designing the electrode spacing is a critical step.
@@
Based on the thrust measurement experiments and the data in Figure 7, we concluded the following: when the spacing is greater than $2.5\text{\,}\mathrm{cm}$, a smaller gap increases the charge density and electric field strength, thereby increasing thrust. However, when the spacing is reduced to around $2.5\text{\,}\mathrm{cm}$, the tip discharge from the positive electrode becomes too intense, creating an overly strong electric field that punctures the aluminum foil.
For the final design, we improved the thruster to the four-ring structure shown in Figure 3 to achieve greater thrust. This version has a net weight of $19.64\text{\,}\mathrm{g}$. With an electrode spacing of $3.0\text{\,}\mathrm{cm}$, the measured maximum thrust was $0.051\text{\,}\mathrm{N}$. This gives the thruster a maximum thrust-to-weight ratio of $2.597\text{\,}\mathrm{N}\text{\,}{\mathrm{kg}}^{-1}$, a significant improvement over the previous dual-ring version.
@@ 3.2.2 Continuously Adjustable Thrust
By commanding throttle via RC, the high-voltage drive is modulated and the thruster produces approximately linear thrust vs. throttle, while the HV-side current changes more than the LV-side current. Figures 8 and 9 show representative relationships.
@@ 4.1 Reference Frames
Since the aerodynamics are simplified to the characteristics of an axisymmetric shape, a new airflow frame is defined: $x_{a}$ is along the freestream direction, $z_{a}$ lies in the $x_{b}x_{a}$ plane and is perpendicular to $x_{a}$, and $y_{a}$ is determined according to the right-hand rule. The angle of attack $\alpha$ is defined as the angle between the airship axis $x_{b}$ and the airflow frame $x_{a}$, and the sideslip angle $\beta$ is defined as the angle between the $x_{b}x_{a}$ plane of the freestream and the airship symmetry plane $x_{b}z_{b}$. Thus, the transformation from the airflow frame to the body frame is Figure  10:
@@ 4.2 External Forces on the Blimp
Study the thrust effects of the thruster in the body frame(Figure 1112). Denote the yaw deflection angle introduced by the servo as $\delta_{y}$, the pitch deflection angle as $\delta_{p}$, the thruster installation position as $s$, and the thruster link length as $l$. Then the thrust force and moment generated by the thruster are:
@@ 5.2 Inner-Loop Stabilization about a Trim
Below, the parameter tuning is illustrated using $\Delta u$ as an example. According to the finite element calculations of the PUB aerodynamic characteristics, we have $C_{D}=0.0071$. As previously mentioned, the mass is $m=0.2978$ kg, and the air density is taken as $\mathfrak{p}=1.205$ kg/m3 (all parameters in SI units). The flight speed is set to 1 m/s. By applying the Laplace transform to the equation with the introduced state feedback, the transfer function model is obtained as shown in Figure 13.
@@
To eliminate the steady-state error of the system in response to a step input, the gain is designed as $k_{u}=0.9828$. A simulation is built in MATLAB/Simulink, and the system step response obtained is shown in Figure 14. By analyzing the response curve, it can be seen that the controller has strong tracking performance and good steady-state characteristics.
@@ 6 Prototype Experiments
Figure . 15 shows the speed measurement experiment of the PUB. By capturing four sets of flight position images at 3-second intervals, the maximum flight speed of the PUB was calculated to be 0.32 m/s. The initial flight altitude was 1.8 m, and after 12 seconds, the altitude decreased to 1.6 m, demonstrating good low-altitude flight stability.
@@
Figure 16 shows the motion capture images for the yaw verification experiment of the PUB. In the experiment, a fixed viewpoint facing along the body axis $Ox_{a}$ was used for recording. From the four motion capture images taken at 3-second intervals in Figure 13, it can be seen that when the rudder on the remote controller was pushed to the right at 0 s, the servo on the two-degree-of-freedom gimbal was able to quickly adjust the yaw angle. The above experiment demonstrates that the remote controller can effectively control both the yaw and pitch angles of the PUB.

### v2

@@ 2.1.1 Blimp Design
As shown in Fig. 1(a), the main body of PUB is an ellipsoidal aluminum-film helium balloon of length $1.97\text{\,}\mathrm{m}$ and maximum diameter $51\text{\,}\mathrm{c}\mathrm{m}$. The volume is approximated by the ellipsoid formula
@@
The thruster layout has undergone several iterations. In the first-generation design (Fig. 1(a)), six identical thrusters are rigidly attached at different positions, requiring simultaneous control of their on/off states and thrust magnitudes to realize six-DOF motion. This resulted in high system mass and control complexity. The second-generation design (Fig. 1(b)) simplifies the layout to a single thruster mounted on a two-DOF platform directly beneath the envelope, enabling thrust vectoring in pitch and yaw and reducing weight while enhancing controllability.
@@ 2.1.2 Two-DOF Gimbal
As shown in Fig. 2, the gimbal uses two RC servos (each $9\text{\,}\mathrm{g}$; total gimbal mass $\approx$33 $3\text{\,}\mathrm{g}$; rated torque $1.6\text{\,}\mathrm{k}\mathrm{g}\,\mathrm{c}\mathrm{m}$) to vector the thruster. The upper stage rotates in the blimp’s oxzoxz symmetry plane to control pitch, with an angular range of $0^{\circ}$$180^{\circ}$. The lower stage is rigidly attached beneath the blimp and rotates about the $z$ axis to control yaw, also within $0^{\circ}$$180^{\circ}$.
@@ 2.2 Plasma Thruster
Fig. 3 shows the thruster: a four-layer concentric ring frame with $9.44\text{\,}\mathrm{m}\mathrm{m}$ spacing between adjacent rings. The positive electrode is $0.1\text{\,}\mathrm{m}\mathrm{m}$ copper wire; the negative electrode is $40\text{\,}\mathrm{m}\mathrm{m}$-wide aluminum foil. The electrode separation is $25\text{\,}\mathrm{m}\mathrm{m}$.
@@ 3.1.2 Principle of Ionic-Wind Propulsion
The effect states that a properly shaped electrode pair immersed in a dielectric and biased at high voltage generates a net force (Fig. 5).
@@
Based on the ion wind effect model from section 3.1.1, we can further simplify and form a thrust model for the ion propulsion device [12] [13]. Fig. 6 illustrates the thrust direction.
@@ 3.2.1 Thrust Measurement
According to the plasma propulsion principles in section 3.1.2, the distance between the positive copper wire and the negative aluminum foil affects the thrust. To find the optimal spacing, we repeatedly changed the gap of the dual-ring thruster and measured the thrust. The experimental process was as follows: The thruster was inverted on a high-precision electronic scale ($0.01\text{\,}\mathrm{g}$ accuracy) with the negative electrode facing up, ensuring it was well-insulated from the scale. After taring the scale, the reading displayed the maximum thrust in grams when the power was turned on. The dual-ring thruster had a net weight of $16.00\text{\,}\mathrm{g}$. We conducted five parallel experiments, decreasing the electrode spacing each time. The thrust measurements are shown in Fig. 7. We found that as the spacing decreased, the thrust gradually increased, reaching a maximum thrust-to-weight ratio of $0.7105\text{\,}\mathrm{N}\text{\,}{\mathrm{kg}}^{-1}$. However, when the gap was reduced to $2.5\text{\,}\mathrm{cm}$, a tip discharge phenomenon occurred, causing the negative aluminum foil to be punctured and burned. This demonstrated that designing the electrode spacing is a critical step.
@@
Based on the thrust measurement experiments and the data in Fig. 7, we concluded the following: when the spacing is greater than $2.5\text{\,}\mathrm{cm}$, a smaller gap increases the charge density and electric field strength, thereby increasing thrust. However, when the spacing is reduced to around $2.5\text{\,}\mathrm{cm}$, the tip discharge from the positive electrode becomes too intense, creating an overly strong electric field that punctures the aluminum foil.
For the final design, we improved the thruster to the four-ring structure shown in Fig. 3 to achieve greater thrust. This version has a net weight of $19.64\text{\,}\mathrm{g}$. With an electrode spacing of $3.0\text{\,}\mathrm{cm}$, the measured maximum thrust was $0.051\text{\,}\mathrm{N}$. This gives the thruster a maximum thrust-to-weight ratio of $2.597\text{\,}\mathrm{N}\text{\,}{\mathrm{kg}}^{-1}$, a significant improvement over the previous dual-ring version.
@@ 3.2.2 Continuously Adjustable Thrust
By commanding throttle via RC, the high-voltage drive is modulated and the thruster produces approximately linear thrust vs. throttle, while the HV-side current changes more than the LV-side current. Fig. 8 and 9 show representative relationships.
@@ 4.1 Reference Frames
Since the aerodynamics are simplified to the characteristics of an axisymmetric shape, a new airflow frame is defined: $x_{a}$ is along the freestream direction, $z_{a}$ lies in the $x_{b}x_{a}$ plane and is perpendicular to $x_{a}$, and $y_{a}$ is determined according to the right-hand rule. The angle of attack $\alpha$ is defined as the angle between the airship axis $x_{b}$ and the airflow frame $x_{a}$, and the sideslip angle $\beta$ is defined as the angle between the $x_{b}x_{a}$ plane of the freestream and the airship symmetry plane $x_{b}z_{b}$. Thus, the transformation from the airflow frame to the body frame is Fig. 10:
@@ 4.2 External Forces on the Blimp
Study the thrust effects of the thruster in the body frame(Fig. 1112). Denote the yaw deflection angle introduced by the servo as $\delta_{y}$, the pitch deflection angle as $\delta_{p}$, the thruster installation position as $s$, and the thruster link length as $l$. Then the thrust force and moment generated by the thruster are:
@@ 5.2 Inner-Loop Stabilization about a Trim
Below, the parameter tuning is illustrated using $\Delta u$ as an example. According to the finite element calculations of the PUB aerodynamic characteristics, we have $C_{D}=0.0071$. As previously mentioned, the mass is $m=0.2978$ kg, and the air density is taken as $\mathfrak{p}=1.205$ kg/m3 (all parameters in SI units). The flight speed is set to 1 m/s. By applying the Laplace transform to the equation with the introduced state feedback, the transfer function model is obtained as shown in Fig. 13.
@@
To eliminate the steady-state error of the system in response to a step input, the gain is designed as $k_{u}=0.9828$. A simulation is built in MATLAB/Simulink, and the system step response obtained is shown in Fig. 14. By analyzing the response curve, it can be seen that the controller has strong tracking performance and good steady-state characteristics.
@@ 6 Prototype Experiments
Fig. 15 shows the speed measurement experiment of the PUB. By capturing four sets of flight position images at 3-second intervals, the maximum flight speed of the PUB was calculated to be 0.32 m/s. The initial flight altitude was 1.8 m, and after 12 seconds, the altitude decreased to 1.6 m, demonstrating good low-altitude flight stability.
@@
Fig. 16 shows the motion capture images for the yaw verification experiment of the PUB. In the experiment, a fixed viewpoint facing along the body axis $Ox_{a}$ was used for recording. From the four motion capture images taken at 3-second intervals in Fig. 13, it can be seen that when the rudder on the remote controller was pushed to the right at 0 s, the servo on the two-degree-of-freedom gimbal was able to quickly adjust the yaw angle. The above experiment demonstrates that the remote controller can effectively control both the yaw and pitch angles of the PUB.